Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $119,600 of manufacturing overhead for an estimated allocation base of $92,000 direct material dollars to be used in production. The company has provided the following data for the just completed year:

Purchase of raw materials $ 137,000
Direct labor cost $ 88,000
Manufacturing overhead costs:
Indirect labor $ 121,900
Property taxes $ 8,100
Depreciation of equipment $ 17,000
Maintenance $ 12,000
Insurance $ 9,900
Rent, building $ 31,000
Beginning Ending
Raw Materials $ 27,000 $ 14,000
Work in Process $ 49,000 $ 38,000
Finished Goods $ 73,000 $ 60,000

Required:

1. Compute the predetermined overhead rate for the year.

2. Compute the amount of underapplied or overapplied overhead for the year.

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $38,000 ending balance in Work in Process includes $8,600 of direct materials. Given this assumption, supply the information missing below:Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 PreComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ---Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 Ass
Answer #1

1) Predetermine overhead rate = 119600/92000 = 130% of direct material used

2) Direct material used = 27000+137000-14000 = 150000

Applied overhead = 150000*130% = 195000

Actual overhead = 199900

Under applied overhead = 199900-195000 = 4900

3) Schedule of cost of goods manufactured

Direct material
Beginning raw material 27000
Raw material purchase 137000
Raw material available for use 164000
Less: Ending raw material -14000
Direct material used 150000
Direct labor 88000
Manufacturing overhead 195000
Total manufacturing cost 433000
Beginning work in process 49000
Total cost of work in process 482000
Less: Ending work in process -38000
Cost of goods manufactured 444000

4) Unadjusted cost of goods sold = 73000+444000-60000 = 457000

5) Calculate following

Direct material 8600
Direct labor 18220
Manufacturing overhead (8600*1.30) 11180
Work in process inventory 38000